A flat circle of radius 20 mm is placed in a uniform electric field of magnitude 20500 m N/C. What us the electric flux through the circle when its face is at 33 degree to the field lines?


1
Expert's answer
2022-02-28T09:59:45-0500

The electric flux through the circle

"\Phi=EA\cos\theta\\=
20500*3.14*(0.02)^2*\cos 57^\circ=14\:\rm N\cdot m^2/C"